# Exploring the Nostr Ecosystem: A Study of Decentralization and Resilience

### v1

@@ Appendix A Appendix
A.1. Sample Relay Information Document
    "contact": "wino@nostr.wine",
    "description": "A paid nostr relay for wine enthusiasts and everyone else.",
    "fees": {
        "admission": [
                "amount": 18888000,
                "unit": "msats"
        "retention": []
    "limitation": {
        "auth_required": false,
        "max_event_tags": 2000,
        "max_limit": 1000,
        "max_message_length": 131072,
        "max_subid_length": 71,
        "max_subscriptions": 50,
        "min_pow_difficulty": 0,
        "payment_required": true
    "name": "nostr.wine",
    "payments_url": "https://nostr.wine/invoices",
    "pubkey": "4918eb332a41b71ba9a74b1dc64276cfff592e55107b93baae38af3520e55975",
    "supported_nips": [1, 2, 4, 9, 11, 12, 15, 16, 20, 22, 28, 33],
    "version": "main"
A.2. Additional Figures for Relay Income
@@
A.3. Algorithms
Algorithm 1 compute the traffics over the following graph $G$.
1:function compute_retrieval_count($P$: Post, $(U,V)$: Following Edge)
2:     $S_{P}^{R}\leftarrow$ set of relays that $P$ is replicated in
3:     $S_{U}^{R}\leftarrow$ set of relays that $U$ is using
4:     return $|S_{P}^{R}\cap S_{U}^{R}|$
5:end function
6:function compute_edge_traffics($(U,V)$: Following Edge)
7:     $S_{V}^{P}\leftarrow$ set of posts transmitted over $(U,V)$
8:     $Count\leftarrow$ sum([compute_retrieval_count($P$, $(U,V)$) for $P\in S_{V}^{P}$])
9:     $Traffics\leftarrow$ sum([compute_retrieval_count($P$, $(U,V)\ \times$ size($P$)) for $P\in S_{V}^{P}$])
10:     return $Count$, $Traffics$
11:end function
12:sum([compute_edge_traffics($(U,V)$) for $(U,V)\in G$.edges])
Algorithm 2 Original implementation of post publication.
1:$P\leftarrow$ the post to be published
2:$L^{R}\leftarrow$ list of relays that the user is using
3:for $R\in L^{R}$ do
4:     publish_post($P$, $R$)
5:end for
Algorithm 3 Modified implementation of post publication with relay selection.
1:$P\leftarrow$ the post to be published
2:$L^{R}\leftarrow$ list of relays that the user is using
3:$N\leftarrow$ the number of relays to be published
4:function relay_selection($L^{R}$: List of Relays, $N$: Number of Selected Relays)
5:     Perform relay selection
6:     $L^{R}_{2}\leftarrow$ list of the selected relays
7:     return $L^{R}_{2}$
8:end function
9:$L^{R}_{2}\leftarrow$ relay_selection($L^{R}$, $N$)
10:for $R\in L^{R}_{2}$ do
11:     publish_post($P$, $R$)
12:end for
Algorithm 4 Original Implementation of Post Retrieval.
1:$L^{R}\leftarrow$ list of relays that the user is using
2:$L^{F}\leftarrow$ list of users that the user is following
3:for $R\in L^{R}$ do
4:     add_subscription($R$, $L^{F}$)
5:end for
Algorithm 5 Modified Implementation of Post Retrieval with Optimistic Retrieval.
1:$N\leftarrow$ number of relays for optimistic retrieval
2:$L^{R}\leftarrow$ list of relays that the user is using
3:$L^{F}\leftarrow$ list of users that the user is following
4:for $R$ in $L^{R}$ do                 % Retrieve ”low-priority” posts with optimistic retrieval
5:     $L_{2}^{F}\leftarrow$ [][][ ]
6:     for $F\in L^{F}$ do
7:         $S_{N}\leftarrow$ set of selected $N$ relays in $L^{R}$ to retrieve the posts of $F$
8:         if $R\in S_{N}$ then
9:              $L_{2}^{F}$.add($F$)
10:         end if
11:     end for
12:     add_low_priority_subscription($R$, $L_{2}^{F}$)
13:end for
14:for $R\in L^{R}$ do                 % Retrieve ”high-priority” posts from all relays in the list
15:     add_high_priority_subscription($R$, $L^{F}$)
16:end for
A.4. Additional Tables for the Improvement of Optimistic Retrieval
Wasted Count Wasted Size
Overall Completeness (%) Wasted (Billion) % of Original Wasted (TiB) % of Original
Original - 218.3 - 143.9 -
$N=2$ 99.0 3.8 1.7 2.5 1.7
$N=3$ 99.2 7.7 3.5 5.1 3.5
$N=4$ 99.3 11.6 5.3 7.7 5.4
$N=5$ 99.4 15.4 7.0 10.2 7.1
Table 2. The post completeness and potentially wasted count and size of posts over the following graph, when adopting optimistic retrieval with N = 2, 3, 4, and 5, with 0% ”high-priority” posts.
Wasted Count Wasted Size
Overall Completeness (%) Wasted (Billion) % of Original Wasted (TiB) % of Original
Original - 218.3 - 143.9 -
$N=2$ 99.0 25.3 11.6 16.7 11.6
$N=3$ 99.2 28.8 13.2 19.0 13.2
$N=4$ 99.3 32.3 14.8 21.3 14.8
$N=5$ 99.4 35.7 16.4 23.6 16.4
Table 3. The post completeness and potentially wasted count and size of posts over the following graph, when adopting optimistic retrieval with N = 2, 3, 4, and 5, with 10% ”high-priority” posts.
Wasted Count Wasted Size
Overall Completeness (%) Wasted (Billion) % of Original Wasted (TiB) % of Original
Original - 218.3 - 143.9 -
$N=2$ 99.0 46.7 21.4 30.8 21.4
$N=3$ 99.2 49.9 22.8 32.9 22.9
$N=4$ 99.3 52.9 24.2 34.9 24.3
$N=5$ 99.4 56.0 25.6 37.0 25.7
Table 4. The post completeness and potentially wasted count and size of posts over the following graph, when adopting optimistic retrieval with N = 2, 3, 4, and 5, with 20% ”high-priority” posts.

### v2

@@ Appendix A Appendix
A.1. Ethical Considerations
The relay dataset only comprises public infrastructure information and data that has been voluntarily disclosed by the relay operator. The event dataset includes both user and post information that are publicly available to anyone. To prevent any mishandling of data, we adhere to established ethical procedures for social media data (Townsend and Wallace, 2016, 2017) and only collect publicly available data. Additionally, we do not analyze specific users or posts, nor do we try to link users to their real personal identities.
Our approach to crawling the relays is designed to be minimally intrusive. Each time we connect to a relay, we gather events from a specific time period rather than retrieving all events in the relay. This method is similar to how a typical user refreshes their timeline in a client application (In fact, our crawler is a modified Nostr client). Consequently, we believe this approach imposes minimal load on the relays, ensuring efficient and respectful data collection.
A.2. Overview of Nostr
Figure 11 offers the screenshots of two Nostr clients. [^2] Screenshot of Iris is retrieved from https://www.alza.cz/nostr-decentralizovana-sit-pro-svobodnejsi-sdileni-obsahu-na-internetu, and screenshot of Damus is retrieved from https://github.com/damus-io/damus.
A.3. Sample Relay Information Document
contact”: wino@nostr.wine”,
description”: A paid nostr relay for wine
enthusiasts and everyone else.”,
fees”: {
admission”: [
amount”: 18888000,
unit”: msats
retention”: []
limitation”: {
auth_required”: false,
max_event_tags”: 2000,
max_limit”: 1000,
max_message_length”: 131072,
max_subid_length”: 71,
max_subscriptions”: 50,
min_pow_difficulty”: 0,
payment_required”: true
name”: nostr.wine”,
payments_url”: https://nostr.wine/invoices”,
pubkey”: ”4918eb332a41b71ba9a74b1dc64276
cfff592e55107b93baae38af3520e55975”,
supported_nips”: [1, 2, 4, 9, 11, 12, 15, 16,
20, 22, 28, 33],
version”: main
A.4. Additional Figures for Relay Income
Figure 12 shows the number of users vs. the estimated 6-month income of free and paid relays. The green line is the estimated 6-month cost of running a relay. The three figures are based on different Bitcoin prices.